Please help me!!!!! I have a hard time with numbers.
Rewrite the equation 4x+4y=20 in slope-intercept form.

Answer by tran3209(100)

The slope-intercept form: y=ax+b
4x+4y=20
4y=-4x+20
y=-4x+20/4
y=-x+5